OpenStudy (chrisplusian):

given Cartesian equations: x=5, and y-4=(z-1)/2 write the vector equation of the line in R^3

OpenStudy (irishboy123):

x is always 5 so work the y,z components ie \(\large \dfrac{y-4}{1}=\dfrac{(z-1)}{2} = t\) and patameterise in t